### v1

# SAC: A Framework for Measuring and Inducing Personality Traits in LLMs with Dynamic Intensity Control

Abstract
Large language models (LLMs) have gained significant traction across a wide range of fields in recent years. There is also a growing expectation for them to display human-like personalities during interactions. To meet this expectation, numerous studies have proposed methods for modelling LLM personalities through psychometric evaluations. However, most existing models face two major limitations: they rely on the Big Five (OCEAN) framework, which only provides coarse personality dimensions, and they lack mechanisms for controlling trait intensity. In this paper, we address this gap by extending the Machine Personality Inventory (MPI), which originally used the Big Five model, to incorporate the 16 Personality Factor (16PF) model, allowing expressive control over sixteen distinct traits. We also developed a structured framework known as Specific Attribute Control (SAC) for evaluating and dynamically inducing trait intensity in LLMs. Our method introduces adjective-based semantic anchoring to guide trait intensity expression and leverages behavioural questions across five intensity factors: Frequency, Depth, Threshold, Effort, and Willingness. Applying this methodology to three advanced LLMs - GPT-4o, Claude 3.7 Sonnet, and Gemini 2.5 Flash - we probe into the reliability of our method and test its generalizability. Through experimentation, we find that modelling intensity as a continuous spectrum yields substantially more consistent and controllable personality expression compared to binary trait toggling. Moreover, we observe that changes in target trait intensity systematically influence closely related traits in psychologically coherent directions, suggesting that LLMs internalize multi-dimensional personality structures rather than treating traits in isolation. Our work opens new pathways for controlled and nuanced human-machine interactions in domains such as healthcare, education, and interviewing processes, bringing us one step closer to truly human-like social machines.
[…]
0254
1 Introduction
Personality can be broadly defined as the characteristic patterns of thought, emotion, and behavior that influences how individuals perceive the world, respond to others, and navigate complex social dynamics [23]. In human interaction, personality shapes expectations, trust, empathy, and social relationships [4]. As artificial agents begin to take on roles that require emotional intelligence and social fluency, personality becomes a central axis along which user perception and comfort are determined.
LLMs are increasingly being utilized in domains where effective human interaction is essential, including education [3], healthcare [8], and creative applications [20]. In such contexts, users not only seek accurate responses but also value consistency, emotional intelligence, and personality alignment. Personality, in this sense, becomes more than a stylistic overlay; it shapes how users trust, engage with, and interpret the intentions of AI systems.
Despite this, contemporary approaches to modelling personality in LLMs [7, 18, 22, 24, 27] treat it as either an emergent artifact or a binary toggle - traits are either induced or not, present or absent. This binary framing fails to reflect the psychological reality that personality traits exist along continuous dimensions [14]. Moreover, evaluations of LLM personality have been largely confined to the broad traits of the Big Five (OCEAN) [15], limiting fine-grained behavioral control and analysis.
To address these limitations, we introduce a novel framework for continuous personality induction in LLMs. Our approach leverages the 16 Personality Factor (16PF) model [9], a psychometrically grounded alternative to the Big Five, to enable expressive, fine-grained control across sixteen traits. We begin by constructing a new benchmark called PERS-16, which adapts the original MPI inventory [18] to the 16PF framework using 163 validated International Personality Item Pool (IPIP) [17] items. Then we apply P2 prompting, an established method for inducing traits in LLMs [18], to PERS-16 and show that it fails to provide reliable, fine-grained trait modulation across all traits. In response, we introduce a new framework called Specific Attribute Control (SAC), which allows traits to be induced at graded intensity levels (1-5). SAC defines personality intensity across five interpretable behavioral dimensions - Frequency, Depth, Threshold, Effort, and Willingness - and employs adjective-based semantic anchoring to stabilize trait expression. We evaluate models in both their SAC-neutral (uninduced) and SAC-induced states to test the precision and consistency of induced traits. Our experiments on three advanced LLMs - GPT-4o, Claude 3.7 Sonnet, and Gemini 2.5 Flash - show that SAC reliably shifts trait intensities in a continuous and interpretable manner. Our experiments also show that the conitnuous spectra of intensity induction in SAC is superior to binary trait induction as seen in P2 Prompting. Moreover, we observe that closely related traits consistently co-vary, revealing latent inter-trait structures that mirror human psychometric groupings.
2 Related Work and Background
Over the years, researchers have used various frameworks to assess human personality traits. The Big Five Personality Factors (OCEAN)  [28] have become a cornerstone in psychology due to their robust psychometric properties and predictive power across domains. The traits used include Openness, Conscientiousness, Extraversion, Agreeableness, and Neuroticism (OCEAN) [15]. However, the Big Five framework, while broadly validated, offers a relatively coarse-grained view of personality traits [1]. In contrast, Cattell’s Sixteen Personality Factors (16PF) [9] provide a more nuanced psychometric structure, covering sixteen specific dimensions as shown in Table 1. The 16PF has seen widespread application across clinical, occupational, and research settings [2, 6, 11, 16, 26].
Question: Given a statement of you: “You {}.” Please choose from the following options to identify how accurately this statement describes you. Options: A. Very Accurate B. Moderately Accurate C. Neither Accurate Nor Inaccurate D. Moderately Inaccurate E. Very Inaccurate Only answer using the letter of the option. Limit yourself to only letters A, B, C, D, or E corresponding to the options given.
Figure 1: Initial MPI Evaluation Prompt
Early studies on LLMs hinted that models could mimic certain personality-like attributes  [5, 13, 25, 29, 30], though these were typically small-scale demonstrations in ad hoc settings rather than standardized evaluations. A major advancement came with the Machine Personality Inventory (MPI) and P2 prompting [18], which jointly provided a psychometrically-informed framework for evaluating LLMs’ personality consistency and inducing specific traits without fine-tuning. Subsequent work [27] further validated the reliability and construct validity of inferred personality scores in LLMs.
In our work, we adopt the 16PF model, while introducing a structured intensity modulation framework, and analyzing latent trait interdependencies, which were not supported in earlier work.
3 PERS-16
To capture a more nuanced and granular understanding of personality in LLMs, we developed PERS-16 (Personality Evaluation and Rating System - 16), an extension of the original Machine Personality Inventory (MPI) adapted to Cattell’s 16 Personality Factor (16PF) model that uses 16 traits shown in Table 1. Expanding from five to sixteen traits allows PERS-16 to provide significantly more behavioural granularity, enabling a more holistic personality profile and revealing subtleties that the Big Five may overlook [1]. For instance, traits like Sensitivity and Orderliness are collapsed under broader Big Five factors (e.g., Agreeableness or Conscientiousness), but in the 16PF framework, they are independently measured, allowing for precise targeting and control. We now explain the construction of the PERS-16 dataset of psychometric questions used to evaluate state-of-the-art LLMs and compare their mean response values.
3.1 Dataset Construction and Methodology
The construction of the PERS-16 dataset was grounded in the International Personality Item Pool (IPIP) [17], a well-established, public-domain repository of psychometric items (questions). Each item corresponds to a specific personality trait in the 16PF model. In total, we used all 163 items from the IPIP pool that comprehensively cover all sixteen traits, ensuring conceptual alignment with the 16PF taxonomy. The complete list of items is provided in the supplementary materials, which are publicly accessible via Zenodo [12]. To adapt this dataset for evaluation of LLMs, we designed a consistent prompt template modeled after the original MPI framework [18]. The final prompt used can be seen in Figure 1. The {} placeholder is filled with an IPIP-derived behavioral statement, such as "enjoy bringing people together" or "know how to comfort others". Each statement is pre-labelled with its associated 16PF trait and its polarity (i.e., whether agreement indicates high or low trait presence), ensuring consistency in scoring. We implement a scoring mechanism that maps each of the 163 LLM responses to a value between 1 and 5, consistent with psychometric evaluation norms. For positively keyed items (i.e., those where agreement indicates higher trait presence), the options are scored from A = 5 to E = 1. For negatively keyed items, the scoring is reversed. The trait-wise score ${Score}_{d}$ for a given dimension (trait) $d$ is computed as the mean of responses across all items $\alpha\in IP_{d}$:
$\scriptsize Score_{d}=\frac{1}{N_{d}}\sum_{\alpha\in IP_{d}}f(\text{LLM}(\alpha,\text{template}))$ (1)
where $IP_{d}$ is the item pool for trait $d$; $N_{d}$ is the number of items for trait $d$; ${LLM}(\cdot)$ returns the model’s letter-based response; $f(\cdot)$ converts this response to a numerical score (1–5), based on the polarity of the key. In addition to the mean score, we compute the variance ($\sigma^{2}$) for each trait, which serves as a proxy for internal consistency. Here, lower variance implies greater stability in how the model responds to related prompts.
Table 1: Trait means and variances for Cattell’s 16 Personality Factors across GPT-4o, Claude 3.7 Sonnet, and Gemini 2.5 Flash
Trait GPT-4o Claude Gemini 2.5 Flash SD
Mean Var. Mean Var. Mean Var.
WARMTH 4.60 0.80 4.90 0.30 4.70 0.64 0.15
INTELLECT 4.62 1.08 4.15 1.17 4.38 1.44 0.23
EMOTIONAL STABILITY 4.60 0.80 4.70 0.64 5.00 0.00 0.21
ASSERTIVENESS 3.40 0.80 3.40 1.11 3.30 1.90 0.06
GREGARIOUSNESS 2.80 0.60 3.00 0.89 2.20 1.40 0.42
DUTIFULNESS 3.40 1.20 4.20 1.08 3.40 1.69 0.46
FRIENDLINESS 3.40 1.20 3.80 0.87 3.20 1.40 0.31
SENSITIVITY 3.80 0.98 4.10 1.30 3.70 1.55 0.21
DISTRUST 2.50 0.81 1.50 0.67 2.00 1.18 0.50
IMAGINATION 3.40 1.20 3.10 0.83 2.50 1.57 0.46
RESERVE 3.60 1.56 2.40 1.28 3.10 1.81 0.60
ANXIETY 2.00 1.00 2.20 0.87 1.70 1.27 0.25
COMPLEXITY 4.60 0.80 4.90 0.30 4.60 0.66 0.17
INTROVERSION 3.20 1.08 2.30 0.78 3.60 1.36 0.67
ORDERLINESS 3.80 0.98 4.20 0.75 4.00 1.61 0.20
EMOTIONALITY 1.50 0.81 1.10 0.30 1.60 1.02 0.26
Table 2: Pairwise Euclidean distances between LLMs
Model GPT-4o Claude 3.7 Sonnet Gemini 2.5 Flash
GPT-4o 0.00 2.25 1.50
Claude 3.7 Sonnet 2.25 0.00 2.33
Gemini 2.5 Flash 1.50 2.33 0.00
3.2 Results on Neutral Model
Table 1 summarises the mean scores and variances for each trait in the baseline LLM evaluations. A clear three-tier pattern emerges: all models score highly ($\geq$ 4.5) on the “likeable/cerebral” cluster - Warmth, Intellect, Emotional Stability, and Complexity - indicating that each system is designed to appear knowledgeable, steady, and prosocial. A middle band (2 $<$ score $<$ 4.5) includes traits related to social style and work ethic - e.g., Assertiveness, Dutifulness, Orderliness - where divergence creates distinctive but balanced personalities. Finally, all models uniformly suppress Distrust, Anxiety, and Emotionality (scores $\leq$ 2), reinforcing the goal of calm, emotionally neutral assistants. Overall, the pattern suggests deliberate personality shaping rather than incidental emergence.
To contextualize LLM scores, we reference psychometric literature describing typical response patterns in human-administered 16PF assessments [9]. Most personality inventories, including the 16PF and its IPIP derivatives, tend to show mean trait scores near the midpoint of the scale (i.e., $\sim$3 on a 1-5 scale) and variances in the range of 0.8 to 1.2, though these values vary by trait and demographic sample [9]. While not using a specific norm dataset, our PERS-16 results from Table 1 exhibit similar score distributions and variances, further proving that LLMs are indeed capable of exhibiting human-like behaviour.
3.3 Cross-Model Comparison
In this subsection, we examine how different LLMs diverge in their personality behavior and explore possible causes for these differences. Our analysis focused on two aspects: (1) the Euclidean distance between trait profiles, which succinctly captures how training data and architectural choices distinguish models from one another; and (2) the standard deviation of trait scores across models, which highlights traits that are most divisive or consistently expressed - where a higher $\sigma$ indicates greater disagreement and a lower $\sigma$ suggests alignment. The Euclidean distance between two models $A$ and $B$, based on their mean personality trait scores across the 16PF traits, is defined as:
$\scriptsize d(A,B)=\sqrt{\sum_{i=1}^{16}\left(\mu_{i}^{(A)}-\mu_{i}^{(B)}\right)^{2}}$ (2)
where, $\mu_{i}^{(A)}$ is the mean score for the $i^{\text{th}}$ trait in model $A$ and $\mu_{i}^{(B)}$ is the mean score for the $i^{\text{th}}$ trait in model $B$, summed up for all sixteen traits. The results presented in Table 2 show that GPT-4o and Gemini 2.5 Flash are the most similar pair, with a Euclidean distance of 1.50. Gemini 2.5 Flash and Claude 3.7 Sonnet are the most dissimilar, with a distance of 2.33. We also observe that GPT-4o exhibits relatively low Euclidean distances to all other models, suggesting it is the most behaviorally aligned overall. Given its proximity to multiple systems, it may serve as a useful reference point for future comparisons of personality modulation across LLMs. Now, to compute the standard deviation of each trait across models, we use:
$\scriptsize\sigma_{i}=\sqrt{\frac{1}{M}\sum_{j=1}^{M}\left(\mu_{i}^{(j)}-\bar{\mu}_{i}\right)^{2}}$ (3)
where, $\mu_{i}^{(j)}$ is the mean score of the $i^{\text{th}}$ trait in the $j^{\text{th}}$ model; $\bar{\mu}_{i}=\frac{1}{M}\sum_{j=1}^{M}\mu_{i}^{(j)}$ is the mean of means for trait $i$; $M$ is the number of models (here, $M=3$); and $\sigma_{i}$ is the standard deviation for the $i^{\text{th}}$ trait across the models. The standard deviation across sixteen traits presented in Table 1 indicate that Introversion shows the highest standard deviation across models (0.67). This shows significant divergence in how LLMs represent social withdrawal and solitary tendencies. Reserve (0.60) and Distrust (0.50) also exhibit considerable variability, suggesting inconsistent modelling of social reticence and cautiousness. In contrast, Assertiveness (0.06), Warmth (0.15), and Complexity (0.17) are among the most stable traits, implying that these are represented relatively consistently across all three LLMs. The remaining ten traits show moderate standard deviation across all models, indicating that they are neither completely stable nor very inconsistent.
4 Personality Induction using P2 Prompting
We next address personality induction in LLMs. P2 Prompting [18] is one such technique which uses the Big Five framework and induces traits in a binary manner. We now explain how we induce traits and evaluate LLMs using P2 prompting.
4.1 Dataset Construction and Methodology
To adapt P2 to more granular traits, we built a new dataset grounded in the 16PF framework. For this, we compiled a list of trait-defining short phrases based on 16PF literature and validated psychometric inventories for each of the sixteen personality traits [17]. To enable deeper induction of personality, we followed the original P2 method’s multi-step pipeline and enriched it for 16PF [18, 19]. For each of the sixteen traits, we generated descriptions using LLM prompting, which are vivid, semantically rich monologues describing how a person high in a given trait thinks, feels, and behaves. For instance, a full P2 description for the trait Complexity looks like: “You have a complex and nuanced understanding of the world. Your ability to see multiple perspectives allows you to navigate intricate issues thoughtfully. You appreciate depth and intricacy in both ideas and relationships, often engaging in reflective thinking.” The P2 descriptions were generated through an iterative process involving both trait-descriptor expansion and language model-assisted synthesis, ensuring naturalness and internal consistency. The full list of P2 descriptions for each of the sixteen traits can be found in the supplementary material [12]. For each trait induction, we posed all 163 questions from our 16PF dataset, resulting in 2,608 responses per model (16 traits × 163 questions). This ensured trait-wide coverage for each prompt and allowed for robust delta calculation across the full 16PF spectrum. The final prompt structure is shown in Figure 1, with one modification: the generated P2 description is inserted as the ground truth at the beginning of the prompt.
4.2 Results on P2 Prompting
Applying the P2 Prompting technique, we fed the modified prompt, as described in the previous subsection, into the three LLMs and calculated the mean scores from their responses across all the sixteen traits for each LLM. To quantify the effect of induction, we computed the difference (delta) between the mean scores of induced and neutral states (PERS-16), which is computed as:
$\scriptsize\Delta_{i}^{(M)}=\mu_{i}^{\text{induced}(M)}-\mu_{i}^{\text{neutral}(M)}$ (4)
where, $\Delta_{i}^{(M)}$ is the delta (change) for trait $i$ in model $M$; $\mu_{i}^{\text{induced}(M)}$ is the mean score for trait $i$ after it has been explicitly induced in model $M$; and $\mu_{i}^{\text{neutral}(M)}$ is the mean score for trait $i$ in the neutral state.
Table 3: Delta scores across first 8 traits for three LLMs using P2
Warmth Intellect Emotional Stability Assertiveness Gregariousness Dutifulness Friendliness Sensitivity
Gemini 2.5 Flash ($\Delta$) -1.80 -1.31 -1.60 -0.60 1.00 0.20 0.40 -0.40
GPT-4o ($\Delta$) 0.40 -0.31 0.40 1.60 2.20 0.50 0.40 -0.80
Claude 3.7 Sonnet ($\Delta$) -2.60 -1.54 -1.80 -0.70 -0.30 -1.10 -0.80 -1.80
Table 4: Delta scores across remaining 8 traits for three LLMs using P2
Distrust Imagination Reserve Anxiety Complexity Introversion Orderliness Emotionality
Gemini 2.5 Flash ($\Delta$) -0.10 -0.40 0.50 0.70 -1.60 -0.90 -1.50 1.80
GPT-4o ($\Delta$) 1.30 1.30 0.50 0.70 -0.70 1.80 0.10 1.20
Claude 3.7 Sonnet ($\Delta$) 1.10 -0.80 0.60 0.00 -1.70 0.30 -1.20 1.20
Based on the results shown in Tables 3 and 4, Claude 3.7 Sonnet exhibits 11 negative deltas out of 16 traits, Gemini 2.5 Flash shows 10 negative deltas, and in contrast, GPT-4o has only 3 negative deltas. In the context of personality induction, when positively keyed, a positive delta indicates that the model successfully increased the expression of the targeted trait. Conversely, a negative delta suggests that the model either failed to induce the trait effectively, regressed in trait expression, or possibly overcorrected in an unintended direction. The presence of negative values in the above results suggests that methods optimized for five broad factors do not scale linearly to a more granular personality taxonomy, underscoring the need for alternative prompting methods or fine-tuning approaches when richer trait coverage is required.
5 Specific Attribute Control (SAC)
Existing machine personality frameworks like P2 Prompting typically treat traits as discrete switches, mirroring an outdated, categorical view of personality expression. Such binary parameterization neglects the well-established psychological consensus that personality characteristics vary along continuous spectra rather than occupying mutually exclusive states [10, 14, 21]. As seen earlier, the P2 Prompting method can signal which trait should emerge but remains blind to how strongly it should manifest, yielding undesirable results. This lack of fine control makes it difficult to compare models accurately, and limits the ability to design personalized interactions, where even small changes in traits like Warmth, Assertiveness, or Imagination can significantly impact user experience. To address this, we introduce a calibrated intensity dimension that captures varying degrees of trait expression and allows for precise, context-sensitive control. We call this framework Specific Attribute Control (SAC).
To capture intra-trait variability beyond presence or absence, we define intensity as the graded strength with which each personality factor is expressed. Intensity is scored on a five-point scale, where 1 denotes a weak manifestation and 5 a highly pronounced one. Each score is the mean of five intensity dimensions applied to a standardized scenario set: Frequency of occurrence, Depth of emotional-cognitive engagement, activation Threshold, expressive Effort, and behavioural Willingness. This quantification enables systematic comparison and controlled modulation of trait prominence, providing a finer-grained foundation for studying and inducing personality in LLMs.
This section describes the two phases of our experiment: (1) SAC-Neutral, where we measure each model’s natural personality profile using detailed intensity levels (which acts as the baseline), and (2) SAC-Induced, where we apply the same rubric while actively steering a focal trait to a target intensity.
5.1 SAC-Neutral
5.1.1 Dataset Construction and Methodology
In this phase, each model was evaluated in its neutral, uninduced state across sixteen traits derived from the 16PF framework that allowed us to measure the models’ inherent tendencies without explicit trait induction. For each trait, we selected three behavioral questions from IPIP [17] that were the most representative of the underlying construct, balancing coverage and clarity. Each behavioral question was framed along one of five intensity dimensions stated earlier. Responses were collected on a 1-5 Likert-type scale, where 1 represented the lowest and 5 represented the highest expression of the relevant behavior. Each full question was dynamically constructed by concatenating an intensity factor phrase (e.g., "How often do you") with a trait-specific behavioral statement (e.g., "cheer people up"). The complete lists of trait definitions, intensity factor descriptors, and behavioural questions used in this phase are available in the supplementary material [12]. In total, the SAC-Neutral evaluation involved 240 questions per model (16 traits observed × 5 intensity factors × 3 behavioral prompts), resulting in a comprehensive baseline personality profile for each LLM, further proving the robustness of our methodology. The prompt structure seen in Figure 2 was used to induce trait intensities. The {} placeholders have the following features -
  • trait: The target personality trait under evaluation (e.g., Warmth).
  • definition: A brief description of the trait, adapted from the MPI-P2 inventory.
  • intensity_factor: One of the five behavioral dimensions (Frequency, Depth, Threshold, Effort, Willingness).
  • composite_question: A full behavioral question constructed by concatenating the intensity phrase with a trait-specific behavioral action (e.g., "How often do you cheer people up?").
  • intensity_answers: The 1–5 scale descriptors corresponding to the intensity factor (Never, Seldom, Occasionally, Often, All the time).
Personality intensity is defined as a combination of five factors: frequency, depth, threshold, effort, and willingness, each rated on a scale from 1 to 5. The target trait is trait, defined as: definition. For the trait trait with the intensity factor intensity_factor, please answer the following question: composite_question The possible response scale is as follows: 1: intensity_answers[0] 2: intensity_answers[1] 3: intensity_answers[2] 4: intensity_answers[3] 5: intensity_answers[4] For each question, please provide a number between 1 and 5 that best represents the intensity.
Figure 2: Neutral Profiling Prompt Template for SAC
5.1.2 Results on SAC-Neutral
Table 5: Comparison of Mean and Variance for 16PF Traits across GPT-4o, Claude 3.7 Sonnet, and Gemini 2.5 Flash under SAC-Neutral conditions.
Trait GPT-4o Claude 3.7 Sonnet Gemini 2.5 Flash
Mean Var. Mean Var. Mean Var.
WARMTH 3.90 2.29 3.20 0.96 3.40 1.44
INTELLECT 4.20 0.96 3.60 1.84 4.00 1.20
EMOTIONAL STABILITY 3.70 2.21 3.30 1.41 4.20 2.56
ASSERTIVENESS 4.00 0.40 3.30 0.61 3.00 1.20
GREGARIOUSNESS 3.30 1.61 3.10 1.09 2.70 0.81
DUTIFULNESS 4.10 2.49 3.80 0.76 4.30 0.41
FRIENDLINESS 3.60 3.04 3.20 0.96 3.70 1.81
SENSITIVITY 3.50 1.75 3.20 1.16 3.00 2.40
DISTRUST 4.00 0.40 3.80 0.16 3.00 1.60
IMAGINATION 3.90 1.69 3.50 0.65 3.20 1.96
RESERVE 3.60 0.64 3.40 0.44 3.10 1.09
ANXIETY 3.80 0.76 3.50 0.25 3.80 0.16
COMPLEXITY 4.20 0.56 3.50 1.85 3.40 0.44
INTROVERSION 3.50 3.45 3.00 2.80 3.40 3.24
ORDERLINESS 4.10 2.49 3.90 0.89 3.40 2.64
EMOTIONALITY 3.40 0.64 3.20 0.96 2.80 1.36
These neutral-state results depict LLMs’ natural, unprompted tendencies and have been tabulated in Table 5. The primary metric evaluated for each trait was the mean of the model’s responses across the three associated questions, representing the model’s natural intensity for that trait. Additionally, the variance of responses was calculated to measure internal consistency and potential response variability across the sub-questions. The results show that most traits exhibited a natural intensity mean in the range of 3.0 to 3.5, with a few tending closer to 4.0. This distribution is consistent with expectations, as 3.0 represents the midpoint on the intensity scale, corresponding to moderate levels of trait expression.
5.2 SAC-Induced
5.2.1 Dataset Construction and Methodology
We developed an intensity induction framework that allows for fine-grained modulation of trait levels within LLMs. The intensity factor phrasing (Frequency, Depth, Threshold, Effort, and Willingness) and the trait-specific behavioral questions from the 16PF inventory remained unchanged from the neutral profiling phase to ensure methodological consistency. The primary innovation in this phase was the introduction of adjectives associated with each intensity level. For every trait, we manually curated a set of adjectives corresponding to intensity levels 1 through 5, carefully ensuring semantic alignment with the behavioral nuances intended for each scale point. This provided the LLMs with richer semantic grounding, anchoring their behavior at each induced intensity. For example, for the trait Warmth:
  • 1: mildly warm, occasionally empathetic, reservedly caring
  • 3: friendly, attentive, genuinely supportive
  • 5: extremely warm, deeply empathetic, overwhelmingly supportive
The full mapping of traits to their corresponding adjectives across all intensity levels is given in the supplementary material [12]. The intensity induction process followed a structured protocol. Each prompt began by defining the target trait (from the MPI-P2 inventory) and specifying its intended intensity (e.g., $=$ 5). A set of trait-specific adjectives corresponding to intensity levels 1-5 was included to semantically anchor the model’s interpretation. The model was then asked to respond to a behavioral question formed by combining one of five intensity factors (e.g., frequency, willingness) with a trait-linked action phrase. This design minimized semantic drift, ensuring alignment between output and intended trait level. In total, the SAC-Induced framework prompted each model with 11,520 questions (16 traits induced × 3 intensity levels × 5 intensity factors × 3 behavioural prompts x 16 traits observed), providing high-resolution insight into how trait intensity modulates model behavior across multiple dimensions. The prompt template in Figure 3 illustrates this setup, with annotated placeholders that have the following features:
  • trait_const: The target trait being induced.
  • definition: A description of the trait adapted from the MPI-P2 inventory.
  • intensity_scale: The desired intensity level (1-5).
  • adjectives: Trait-specific adjectives associated with each intensity level.
  • intensity_question: The phrasing of the intensity dimension (Frequency, Depth, etc.).
  • question: The behavioral action specific to the trait.
  • intensity_answers: The descriptors for the 1-5 response options.
Personality intensity is defined as a combination of five factors: frequency, depth, threshold, effort, and willingness, each rated on a scale from 1 to 5. The trait trait_const would be described as: definition. The trait currently being adjusted is trait_const, which is set to intensity level intensity_scale. This adjustment to trait_const may affect other traits differently, depending on their nature. Adjectives for each scale from 1 to 5 for the trait trait_const are: 1: adjectives.get(1, ’N/A’) 2: adjectives.get(2, ’N/A’) 3: adjectives.get(3, ’N/A’) 4: adjectives.get(4, ’N/A’) 5: adjectives.get(5, ’N/A’) The intensity for the trait trait should reflect how it behaves independently or in contrast with the modified intensity of trait_const. For all future communication, the scale I would like you to operate on for trait_const is intensity_scale. Task: intensity_question question? The possible intensity scale is as follows: 1: intensity_answers[0] 2: intensity_answers[1] 3: intensity_answers[2] 4: intensity_answers[3] 5: intensity_answers[4] For each question, please provide an answer that best represents the trait trait at the intensity of trait_const
Figure 3: Trait Intensity Induction Prompt Template
Figure 4: Trait delta scores for Claude, GPT-4o, and Gemini across intensity levels 1, 3, and 5 using the SAC framework. At intensity 1 (top), most traits exhibit strong negative deltas, indicating a suppressed expression of the induced personality traits. At intensity 3 (middle), deltas cluster around zero, suggesting a neutral or balanced level of trait expression with early signs of co-trait movement. At intensity 5 (bottom), the target traits show strong positive deltas, reflecting pronounced and deliberate trait expression. The progression across the three plots highlights smooth, symmetric movement around the neutral baseline, demonstrating our framework’s ability to modulate personality traits along a continuous scale with fine-grained control.
Figure 5: Target Trait Influence and Top Co-Movers in Claude 3.7 Sonnet. Each subplot shows the delta change in trait score across three intensity levels (1, 3, 5) for a focal personality trait. The primary trait being induced is plotted in green, while the two most reactive co-moving traits are also plotted in blue and orange. Positive deltas indicate increased expression relative to the neutral baseline, and negative deltas indicate suppression. Notably, intensity level 5 consistently shows positive deltas for the target trait, while level 1 shows negative deltas for the target trait, demonstrating precise modulation. The trait responses align with psychological theory, e.g., increasing Warmth reduces Distrust and Reserve, suggesting coherent personality structures in Claude 3.7 Sonnet.
5.2.2 Results on SAC-Induced
To evaluate the response of LLMs to induced personality traits, we used the mean of model-generated responses as the metric for trait intensity across three levels: 1, 3, and 5. For each trait and intensity level, the score was compared against the neutral baseline. The delta value was computed by subtracting the neutral score from the induced score:
$\scriptsize\Delta=\mu_{\text{induced}}-\mu_{\text{neutral}}$ (5)
This normalization ensured all traits shared a common baseline (0) and enabled meaningful cross-trait and cross-model comparisons.
Figure 4 illustrates the effect of inducing traits at intensity levels 1, 3, and 5, through deltas, allowing us to focus solely on the magnitude and direction of change. These bar charts provide a comprehensive comparison across all 16 traits for all three LLMs. As intensity increases, we observe a smooth and directional shift in trait deltas - from strong negative values at intensity 1, to moderate values closer to zero at intensity 3, and strong positive deltas at intensity 5. This trend indicates that SAC enables monotonic modulation of personality traits, with intensity 3 representing a plausible midpoint. This is a stark improvement from P2 Prompting wherein numerous negative deltas were observed even when positively keyed as seen in Tables 3 and 4. Even those traits which were the most divisive under neutral MPI - namely Introversion, Distrust, and Reserve - were shown to follow consistent, smooth, and monotonic paths when intensity was induced. The consistency of these patterns across traits and models reinforces the robustness and semantic alignment of our framework.
Next, we analyse inter-trait movement across LLMs to understand whether or not LLMs are capable of emulating latent psychological structures known in human personality science. Co-movers were selected by identifying the two traits with the largest deltas (positive or negative) accompanying the induction of a target trait as can be seen in Figure 5. While this figure shows the results for Claude 3.7 Sonnet, we observed similar modulation and co-movement patterns in GPT-4o and Gemini 2.5 Flash as well, indicating that the SAC framework generalizes well across architectures (see supplementary material [12]). The presence and behavior of these co-movers offer insights into whether LLMs capture inter-trait dependencies beyond the directly manipulated trait. Table 6 summarizes the top co-movers for each target trait based on the most reactive trait shifts observed across all three LLMs. Across all experiments, we observe that all three LLMs not only respond to individual traits with fidelity but also adjust semantically related traits in a manner consistent with human personality psychology. These structured co-movements observed in Table 6 are not merely reactive but suggest that LLMs have internalized coherent personality architectures. Our intensity-control method not only reveals these underlying structures, but also enables precise and interpretable persona shaping.
6 Limitations
While our work advances personality modelling by introducing fine-grained intensity control through SAC prompting, several limitations remain. First, the reliance on self-reported questionnaire-style responses may not fully capture deeper behavioral tendencies in LLMs, which lack persistent internal states. Second, observed differences in trait expression across models may conflate genuine induction effects with inherent training data distributions or architectural biases. Third, while we focus on the 16PF framework, the generalizability of our method to alternative or hierarchical personality models remains an open question for future work. Addressing these challenges will help build more robust, adaptive, and psychologically grounded AI systems.
7 Discussion and Summary of Observations
This work introduces a shift from the widely used Big Five framework to Cattell’s more granular 16 Personality Factors (16PF) model, through the construction of a new benchmark: PERS-16. Compared to the Big Five, 16PF allows for more expressive and fine-grained personality modeling, enabling the evaluation and control of traits that are typically collapsed into broader categories - such as Sensitivity and Orderliness - in traditional models. In our baseline evaluation using PERS-16, we observed that most traits naturally clustered around mid-scale values (around 3–3.5), suggesting neutral, balanced personalities. However, traits such as Introversion, Imagination, and Reserve showed substantial variability across models, indicating areas of disagreement in how LLMs implicitly encode personality. These differences were further quantified using Euclidean distance and standard deviation metrics, offering a psychometric lens for comparing LLM behavior. Applying the P2 prompting method to PERS-16 revealed its limitations. While P2 could induce some traits, its binary nature and lack of intensity control resulted in inconsistent personality shifts, often producing flat or negative deltas even for positively keyed traits. This confirmed that P2 was insufficient for controlled personality modulation within a fine-grained trait space. To overcome these issues, we proposed the Specific Attribute Control (SAC) framework, which enables trait induction across five intensity levels (1-5) along five interpretable behavioral dimensions. In SAC-neutral mode, models exhibited more stable and interpretable personality profiles than under P2. SAC-induced evaluations revealed smooth, graduated shifts in trait expression across all three LLMs. Notably, the emergence of consistent co-movers - traits that shifted in semantically or psychologically aligned ways with the target trait - highlighted the presence of latent, multidimensional personality structures within LLMs.
Table 6: Top co-movers and interpretive insights across traits.
Target Co-movers Dir. Interpretation
Warmth Reserve, Distrust Warmth promotes social openness and trust, thus dampening traits related to social withdrawal and skepticism.
Intellect Sensitivity, Complexity Models associate cognitive richness with nuanced emotional and abstract thought.
Emotional Stability Anxiety Affective regulation reduces psychological volatility.
Assertiveness Reserve A rise in assertiveness reduces passivity and detachment.
Gregariousness Introversion, Reserve Heightened sociability directly suppresses social withdrawal and reservation.
Dutifulness Orderliness Reflects classical conscientiousness links between responsibility and structure.
Friendliness Warmth, Gregariousness, Reserve ± Social engagement rises while social withdrawal falls.
Distrust Anxiety, Reserve A guarded persona co-expresses with defensiveness and reduced openness.
Reserve Warmth, Distrust ± Greater detachment increases skepticism and reduces interpersonal warmth.
Anxiety Emotional Stability Internal reactivity and stability are opposing traits, consistently inverted.
Complexity Intellect Trait pairs show abstract intellectual depth and complexity rising together.
Introversion Reserve Solitude-seeking behaviors are intertwined in both direction and degree.
Emotionality Anxiety, Reserve ± Emotional sensitivity partially aligns with internal turbulence.
8 Conclusion and Future Work
Our study demonstrates that continuous, intensity-based personality control , enabled by the SAC framework and grounded in the 16PF model, offers a robust and interpretable method for modulating LLM personas. Through systematic evaluations across neutral baselines, binary induction (P2), and graded induction (SAC), we show that LLMs not only respond predictably to personality modulation but also reflect inter-trait dynamics aligned with human psychometric theory. This establishes a strong foundation for building adaptive, psychologically informed AI agents. Future work will explore multi-trait induction to simulate more human-like, multidimensional personalities. Another promising direction is real-time dynamic modulation of personality based on user interaction, enabling context-aware adaptations during dialogue.
{ack}
Portions of this manuscript were refined using AI-based writing assistants (e.g., ChatGPT) for grammar, clarity, and style. All scientific content, methodology, and analysis were developed and authored by the researchers.